# STEPC: A Pixel-wise Nonuniformity Correction Framework for Photon-Counting CT in Multi-material Imaging Scenarios

### v1

@@ 6 DISCUSSION
In this study, we proposed the Signal-to-Uniformity Error Polynomial Calibration (STEPC) method to address detector response nonuniformity in photon-counting CT systems. It begins by generating ideal flat-field projections using a 2D second-order polynomial fit, eliminating the need for precise physical thickness measurements of calibration phantoms. An empirical polynomial model is then built on the residuals across all energy thresholds. This approach overcomes the limitation of requiring more energy thresholds for multi-material calibration, enabling accurate prediction and correction of nonuniformity errors under complex incident spectra.
@@
Detector nonuniformity is dependent on the incident X-ray spectra as indicated in equation 12. Previous studies have used slab phantoms with varying thicknesses of PMMA (to simulate soft tissue) and aluminum (to simulate bone) to perform single or dual-material calibration. However, no calibration strategies have yet been developed specifically for contrast agents. We found that the choice of calibration materials is critical. As shown in Figure 9, STC calibrated only with PMMA performs well for PMMA but poorly for CaCl2 phantom. Similarly, PMDC, calibrated with both PMMA and aluminum, achieves better correction for both PMMA, CaCl2 phantoms and mouse head, but its performance degrades in the presence of contrast agents, as seen in Figure 10 and Figure 11. Notably, nonlinear spectral effects are more prominent at lower energy bins due to increased beam hardening and material-specific absorption variations. This significantly challenges calibration models that assume linearity, such as ATC, which exhibits more severe ring artifacts in low-energy images, as shown in Figure 9-11. To address these challenges, we designed a dedicated slab phantom containing an iodinated contrast solution to calibrate contrast-enhanced objects. However, most photon-counting detectors are currently limited to two energy thresholds, making material decomposition methods like PMDC unsuitable for three-material decomposition. The proposed STEPC method overcomes this limitation by directly modeling and predicting nonuniformity errors without relying on material decomposition and additional energy thresholds. By employing a nonlinear multi-energy polynomial model, STEPC effectively captures spectral nonlinearity and enables more accurate and robust correction in complex multi-material imaging scenarios.
@@
On the other hand, the proposed STEPC method can be flexibly adapted to different imaging scenarios. For non-contrast-enhanced mouse scans, using only PMMA and aluminum in calibration is sufficient to approximate the possible incident spectra, as shown in Figure 1. For contrast-enhanced scans, only an additional slab phantom containing the contrast agent needs to be included, no changes to the model itself are required. STEPC also demonstrates strong robustness, successfully suppressing ring artifacts across single-material, dual-material, and even triple-material phantoms with iodinated contrast, as shown in Figure 9-10. In vivo results also confirm its superior performance in both non-contrast mouse head and contrast-enhanced kidney imaging (Figure 11). Moreover, STEPC offers greater implementation flexibility and low computational cost. Unlike thickness-based correction methods such as STC and PMDC, which require precise phantom positioning and accurate thickness measurements, STEPC uses a 2D second-order polynomial to estimate the ideal projection surface, simplifying the calibration process. Meanwhile, As demonstrated in the Figure 12 and Table 3, a second-order polynomial is sufficient to achieve high correction accuracy, significantly reducing computational demands.
Despite its advantages, this work has several limitations and directions for future research:
@@
    Extension to Complex Scenarios: This study focuses on dual-threshold detectors. As future photon-counting CT systems support more energy thresholds, STEPC could be extended to higher-dimensional spectral spaces to better handle multi-contrast-agent cases. It is also necessary to assess the generalizability of iodine-based calibration to other contrast agents (e.g., gadolinium or barium).
@@
7 Conclusion
In this study, we proposed the STEPC method to address detector nonuniformity in photon-counting CT. STEPC enables effective correction for multi-material objects even with only two energy thresholds. Experimental results demonstrate that, compared to existing single- and multi-energy correction methods, STEPC more effectively leverages spectral information and shows superior robustness and generalizability, particularly in contrast-enhanced imaging, where it achieves better artifact suppression. Overall, STEPC offers greater operational flexibility and adaptability for complex multi-material scenarios. Therefore, these qualities position it as a promising general-purpose calibration framework for photon-counting CT systems.
[…]

### v2

@@ VI DISCUSSION
In this study, we proposed the Signal-to-Nonuniformity Error Polynomial Calibration (STEPC) method to address detector response nonuniformity in photon-counting CT systems. It begins by generating ideal flat-field projections using a 2D second-order polynomial fit. An empirical polynomial model is then built on the residuals across all energy thresholds. This approach overcomes the limitation of requiring more energy thresholds for multi-material calibration, enabling accurate prediction and correction of nonuniformity errors under complex incident spectra.
@@
Detector nonuniformity is dependent on the incident X-ray spectra as indicated in Eq. 12. Previous studies have used slab phantoms with varying thicknesses of PMMA (to simulate soft tissue) and aluminum (to simulate bone) to perform single or dual-material calibration. However, no calibration strategies have yet been developed specifically for contrast agents. We found that the choice of calibration materials is critical. As shown in Fig. 10, STC calibrated only with PMMA performs well for PMMA but poorly for CaCl2 phantom. Similarly, PMDC, calibrated with both PMMA and aluminum, achieves better correction for both PMMA, CaCl2 phantoms and mouse head, but its performance degrades in the presence of contrast agents, as seen in Fig. 11 and Fig. 12. Notably, nonlinear spectral effects are more prominent at lower energy bins due to increased beam hardening and material-specific absorption variations. This significantly challenges calibration models that assume linearity, such as ATC, which exhibits more severe ring artifacts in low-energy images, as shown in Fig. 10 - 12. To address these challenges, we designed a dedicated slab phantom containing an iodinated contrast solution to calibrate contrast-enhanced objects. However, most photon-counting detectors are currently limited to two energy thresholds, making material decomposition methods like PMDC unsuitable for three-material decomposition in the projection domain. The proposed STEPC method overcomes this limitation by directly modeling and predicting nonuniformity errors without relying on material decomposition and additional energy thresholds. By employing a nonlinear multi-energy polynomial model, STEPC effectively captures spectral nonlinearity and enables more accurate and robust correction in complex multi-material imaging scenarios.
@@
On the other hand, the proposed STEPC method can be flexibly adapted to different imaging scenarios. For non-contrast-enhanced mouse scans, using only PMMA and aluminum in calibration is sufficient to approximate the possible incident spectra, as shown in Fig. 1. For contrast-enhanced scans, only an additional slab phantom containing the contrast agent needs to be included, no changes to the model itself are required. STEPC also demonstrates strong robustness, successfully suppressing ring artifacts across single-material, dual-material, and even triple-material phantoms with iodinated contrast, as shown in Fig. 10 and Fig. 11. In vivo results also confirm its superior performance in both non-contrast mouse head and contrast-enhanced kidney imaging (Fig. 12). Moreover, STEPC can be readily extended for beam hardening correction, enabling simultaneous suppression of both ring and beam-hardening artifacts, as demonstrated in Fig. 13. The beam-hardening correction module can also operate independently or be integrated with other calibration methods such as FF, STC, or ATC (Fig. 14), making the framework applicable to more complex imaging conditions. Regarding the quantitative impact on material decomposition, we evaluated the accuracy of calcium and iodine quantification. Although STEPC did not achieve the best precision for calcium, the quantitative error remained small and within acceptable limits, while iodine quantification was the most accurate among all methods.
Despite its advantages, STEPC may be susceptible to overfitting due to the use of higher-order polynomial models. Overfitting mainly arises from noise and material variability. To suppress noise-induced overfitting, sufficient frame averaging is required during calibration; in this work, 600 frames were acquired for each slab combination and averaged for model fitting. Material overfitting depends on the selection and combinations of calibration materials. For biological imaging, appropriate combinations of PMMA and aluminum slabs with different thicknesses should be selected to effectively cover typical attenuation ranges, as illustrated in Fig. 1. When the number of calibration combinations needs to be reduced, overfitting risk can be mitigated by selecting thickness combinations that yield a broad and well-distributed range of attenuation values. However, for materials not included in calibration, such as iodine, whose attenuation characteristics differ substantially from those of PMMA and aluminum, correction performance may degrade, particularly for higher-order models, as shown in Table V and Fig. 15. In such cases, incorporating additional calibration materials can improve robustness for more diverse imaging scenarios. Empirically, second-order model already achieves effective correction, with only marginal improvement from higher-order models. Nevertheless, for higher-order polynomials or limited calibration data, overfitting may still occur, and additional strategies such as parameter regularization, cross-validation on a subset of calibration slabs, or averaging over more frames could be considered in future implementations. Consequently, limiting the polynomial order to no higher than third order is recommended to balance correction accuracy, robustness, and computational efficiency.
All experiments in this study were conducted on a single custom Micro-PCCT system. In principle, however, STEPC can be readily extended to other Micro-PCCT systems and even to clinical PCCT platforms. Different systems may exhibit distinct detector designs and spectral characteristics, and additional system-specific corrections may be required prior to applying STEPC. For example, clinical PCCT systems typically operate under higher X-ray flux conditions, where pile-up effects become non-negligible and should be corrected before nonuniformity calibration. Changes in energy threshold settings, tube voltage, or filtration conditions modify the effective X-ray spectrum and detector response, and therefore recalibration is required to maintain correction accuracy, which is standard practice for PCCT calibration methods. Regarding robustness, the ideal reference projections in STEPC are generated via second-order 2D polynomial surface and multi-frame averaging across all detector pixels, which mitigates statistical noise and avoids introducing spatial bias. As demonstrated in Figs. 1012, STEPC does not amplify noise or induce center-to-edge artifacts under the studied acquisition conditions. However, for more complex system configurations, such as those employing bowtie filter or operating in extremely low-count regimes, noise propagation and model robustness may require additional consideration. In such cases, additional regularization strategies or increased calibration averaging may be beneficial to preserve correction stability. For other contrast agents, such as gadolinium, STEPC can in principle be applied following the same calibration procedure used for iodine-based imaging. In addition, when a long-time interval exists between calibration and data acquisition, detector drift and small variations in the X-ray spectrum may degrade correction performance, which is a common issue in photon-counting CT systems [34]. Periodic recalibration, for example every few months, is therefore recommended to maintain correction stability.
Future work may focus on the following directions:
@@
    Extension to Complex Scenarios: This study focuses on dual-threshold Micro-PCCT systems. As next-generation photon-counting detectors support more energy thresholds, STEPC could be generalized to higher-dimensional spectral spaces to better handle multi-contrast-agent imaging. It is also necessary to assess the generalizability of iodine-based calibration to other contrast agents (e.g., gadolinium or barium).
@@
VII Conclusion
In this study, we proposed the STEPC framework to address measurement nonuniformity of PCDs in Micro-PCCT systems under multi-material imaging conditions. STEPC treats detector nonuniformity as a measurable system error and employs a multi-energy polynomial formulation to implicitly incorporate both non-ideal detector response and material-dependent attenuation effects, enabling accurate prediction and correction of nonuniformity errors in the projection domain. By calibrating the model using multi-material slab phantoms, the proposed framework achieves effective correction for complex multi-material objects, even when only two energy thresholds are available. Experimental results demonstrate that, compared with existing single- and multi-energy correction approaches, STEPC more effectively exploits spectral measurement information, leading to improved projection uniformity and substantial suppression of ring artifacts in both multi-material phantom experiments and mouse imaging with and without contrast agents. In addition, the framework can be naturally extended to compensate for beam hardening effects. Overall, STEPC provides a flexible and robust calibration strategy for modeling and correcting PCD measurement nonuniformity, making it a potentially general-purpose calibration framework for Micro-PCCT systems.
Appendix A Derivation of Eq. (12)
The pixel-wise nonuniformity error can be expressed as:
$N_{k,j}^{\mathrm{error}}=N_{k,j}-N_{k,j}^{\mathrm{ideal}}$ (32)
According to Eq. (11), the ideal photon counts can be rewritten as
$\begin{split}&N_{k,j}^{\mathrm{ideal}}=\\ &\int_{E_{k}-\Delta E_{k,j}}^{E_{\max}}\int_{0}^{E_{\max}}S(E)e^{-\int_{l}{\mu(E,\vec{x})dl}}R_{j}^{\mathrm{ideal}}(E^{\prime},E)\,dE\,dE^{\prime}\\ &-\int_{E_{k}-\Delta E_{k,j}}^{E_{k}}\int_{0}^{E_{\max}}S(E)e^{-\int_{l}{\mu(E,\vec{x})dl}}R_{j}^{\mathrm{ideal}}(E^{\prime},E)\,dE\,dE^{\prime}\end{split}$ (33)
Substituting Eq. (10) and Eq. (33) into Eq. (32), the nonuniformity error can be expressed as
$\begin{split}&N_{k,j}^{\mathrm{error}}=\\ &\int_{E_{k}-\Delta E_{k,j}}^{E_{\max}}\int_{0}^{E_{\max}}S(E)e^{-\int_{l}{\mu(E,\vec{x})dl}}\Delta R_{j}(E^{\prime},E)\,dE\,dE^{\prime}\\ &+\int_{E_{k}-\Delta E_{k,j}}^{E_{k}}\int_{0}^{E_{\max}}S(E)e^{-\int_{l}{\mu(E,\vec{x})dl}}R_{j}^{\mathrm{ideal}}(E^{\prime},E)\,dE\,dE^{\prime}\end{split}$ (34)
In practical photon-counting detectors, the threshold offset $\Delta E_{k,j}$ is typically small compared with the energy bin width. Under this assumption, the above expression can be approximated as:
$\begin{split}&N_{k,j}^{\mathrm{error}}\approx\\ &\int_{E_{k}}^{E_{\max}}\int_{0}^{E_{\max}}S(E)e^{-\int_{l}{\mu(E,\vec{x})dl}}\Delta R_{j}(E^{\prime},E)\,dE\,dE^{\prime}\\ &+\Delta E_{k,j}\int_{0}^{E_{\max}}S(E)\,e^{-\int_{l}\mu(E,\vec{x})\,dl}\,R_{j}^{\mathrm{ideal}}(E_{k},E)\,dE.\end{split}$ (35)
Eq (35) corresponds to the approximate nonuniformity error model given in Eq. (12) of the main text.